OpenStudy (anonymous):

What's the intercept form for y=x ^2-2x-3 and y=-2x^2-4x+3

OpenStudy (anonymous):

Intercept? there are lots of places (lines) they can intercept buddy. those 2 are already in y intercept form. If you want them in x intercept then ust the quadratic formula is one of the many ways.

OpenStudy (ranga):

Intercept form for a quadratic is expressed as: y = a(x-p)(x-q). Essentially you need to factor the quadratic expression. y = x ^2 - 2x - 3 y = x^2 - 3x + x - 3 y = x(x - 3) + 1(x-3) y = (x-3)(x+1) Now you can try to factor y=-2x^2-4x+3

OpenStudy (anonymous):

I got (x-9) (x-2), but feel like it isn't correct? @ranga

OpenStudy (ranga):

The second one does not factor easily. So you use the quadratic formula and find the roots: { -b +/- sqrt( b^2 - 4ac) } / 2a

OpenStudy (anonymous):

For the roots I got x= 1/2 (-2+√ 10), -1/2 (2+√ 10) @ranga

OpenStudy (ranga):

You can perhaps put it in decimal.

OpenStudy (anonymous):

I get -3.16 @ranga